# URLOST: Unsupervised Representation Learning without Stationarity or Topology

### v1

@@ Appendix A Appendix
A.1 Spectral clustering algorithm
@@
Let $A_{ij}=I(X_{i};X_{j})$ be the affinity matrix, $p(i)$ be the density function defined in 4. We follow the steps from [38] to perform spectral clustering with a modification to adjust the density:
@@
    Define $D$ to be the diagonal matrix whose ($i$,$i$)-element is the sum of $A$’s $i$-th row, $P$ be the identity matrix where $P_{ii}=p(i)$. Construct the matrix $L=P^{\frac{1}{2}}D^{-\frac{1}{2}}AD^{-\frac{1}{2}}P^{\frac{1}{2}}.$
@@
Some other interpretation of spectral embedding allows one to design a specific clustering algorithm in step 4. For example, [56] interprets the eigenvector problem in 1 as a relaxed continuous version of K-way normalized cuts problem, where they only allow $X$ to be binary, i.e. $X\in\{0,1\}^{N\times K}$. This is an NP-hard problem. Allowing $X$ to take on real value relaxed this problem but created a degeneracy solution. Given a solution $X^{*}$ and $Z=D^{-\frac{1}{2}}X^{*}$, for any orthonormal matrix $R$, $RZ$ is another solution to the optimization problem 1. Thus, [56] designed an algorithm to find the optimal orthonormal matrix $R$ that converts $X^{*}$ to discrete value in $\{0,1\}^{N\times K}$. From our experiment, [56] is more consistent than K-means and other clustering algorithms, so we stick to using it for our model.
A.2 Data synthesize process
We followed the retina sampling approach described in [8] to achieve foveated imaging. Specifically, each retina ganglion cell is represented using a Gaussian kernel. The kernel is parameterized by its center, denoted as $\vec{x}_{i}$, and its scalar variance, $\sigma_{i}^{\prime 2}$, i.e. $\mathcal{N}(\vec{x}_{i},\sigma_{i}^{\prime 2}\mathbf{I})$, which is illustrated in Figure 5.A. The response of each cell, denoted as $G[i]$, is computed by the dot product between the pixel value and the corresponding discrete Gaussian kernel. This can be formulated as:
@@ A.2 Data synthesize process
Table 4: Evaluation on foveated CIFAR-10 with varying hyperparameter for density function. For each set of values of $\alpha$ and $\beta$, we perform density adjusted spectral clustering and run URLOST with the corresponding cluster. The evaluation of each trial is provided in the table.
@@
A.3 Density adjusted spectral clustering on foveated CIFAR10 dataset
We provide further intuition and visualization on why density adjusted spectral clustering allows the model to learn a better representation on the foveated CIFAR-10 dataset.
As shown in Figure 5, the kernel at the center is much smaller in size than the kernel in the peripheral. This makes the kernel at the center more accurate but smaller, which means it summarizes less information. Spectral clustering with constant density will make each cluster have a similar number of elements in them. Since the kernel in the center is smaller, the cluster in the center will be visually smaller, than the cluster in the peripheral. The effect is shown in Figure 6. Moreover, since we’re upsampling an already low-resolution image (CIFAR-10 image), even though the kernel at the center is more accurate, we’re not getting more information. There, to make sure each cluster has similar information, the clusters in the center need to have more elements than the clusters in the peripheral. In order to make the clusters at the center have more elements, we need to weight the clusters in the center more with the density function. Since the sampling kernels at the center have small eccentricity and are more correlated to their neighbor, increasing $\alpha$ and $\beta$ will make sampling kernels at the center have higher density, which makes the cluster at the center larger. This is why URLOST with density adjusted spectral clustering performs better than URLOST with constant density spectral clustering, which is shown in Table 4. Meanwhile, setting $\alpha$ and $\beta$ too large will also hurt the model’s performance because it creates clusters that are too unbalanced.
Figure 6: Effect of density adjusted clustering. Eccentricity-based sampling lattice. The center of the sampling lattice has more pixels which means higher resolution compared to the peripheral. (A) Result of density adjusted spectral clustering ($\alpha=0.5,\beta=2$). Clusters in the center have more elements than clusters in the peripheral. But clusters look more visually similar in size than B. (B) Result of uniform density spectral clustering ($\alpha=0,\beta=0$). Each cluster has a similar number of elements in them but the clusters in the center are much smaller than the clusters in the periphery.
A.4 self-organizing layer learns inverse permutation
@@ A.3 Density adjusted spectral clustering on foveated CIFAR10 dataset
A.5 Visualizing the weight of self-organizing

### v2

@@ Appendix A Appendix
A.1 Motivation of density adjusted spectral clustering
Using the terminologies in functional analysis, the mutual information graph defined in section 2.2 corresponds to a compact Riemannian manifold $\mathcal{M}$ and the Laplacian matrix $L$ is a discrete analogous to the Laplace Beltrami operator $\mathcal{L}$ on $\mathcal{M}$. Minimizing the spectral embedding objective $tr(YLY^{T})$ directly corresponds to the following optimization problem in function space:
[…]
where $f(x):\mathcal{M}\to[0,1]$ is the normalized signal defined on $\mathcal{M}$. We particularly want to write out the continuous form of spectral embedding so we can adapt it to non-stationary signals. To do so, we assume the measure $\lambda$ is absolutely continuous with respect to standard measure $\mu$. By apply the Radon-Nikodym derivative to equation  5, we get:
[…]
where the quantity $\frac{d\lambda}{d\mu}$ is called the Radon-Nikodym, which is some form of density function. Let $p(x)=\frac{d\lambda}{d\mu}$, we can rewrite the optimization problem as the following:
[…]
The density function $p(x)$ on the manifold is analogous to the density adjustment matrix in equation 1. Standard approaches in equation 5 assume that nodes are uniformly distributed on the manifold, thereby treating $p(x)$ as a constant and excluding it from the optimization process. However, this assumption does not hold in our case involving non-stationary signals. Our work introduces a variable density function $p(x)$ for each signal, making it a pivotal component in building good representations for non-stationary signals. This component is referred to as Density Adjusted Spectral Clustering. Empirical evidence supporting this design is provided through visualization and ablation studies in the experimental section.
A.2 Spectral clustering algorithm
@@
Let $A_{ij}=I(S_{i};S_{j})$ be the affinity matrix, and let the density adjustment matrix be $P$ defined in 2.2. Correlation is used instead of mutual information when the dimension is really high since computing mutual information is expensive. We follow the steps from [51] to perform spectral clustering with a modification to adjust the density:
@@
    Define $D$ to be the diagonal matrix whose ($i$,$i$)-element is the sum of $A$’s $i$-th row. Construct the matrix $L=P^{\frac{1}{2}}D^{-\frac{1}{2}}AD^{-\frac{1}{2}}P^{\frac{1}{2}}.$
@@
Some other interpretation of spectral embedding allows one to design a specific clustering algorithm in step 4. For example, [72] interprets the eigenvector problem in 6 as a relaxed continuous version of K-way normalized cuts problem, where they only allow $X$ to be binary, i.e. $X\in\{0,1\}^{N\times K}$. This is an NP-hard problem. Allowing $X$ to take on real value relaxed this problem but created a degeneracy solution. Given a solution $X^{*}$ and $Z=D^{-\frac{1}{2}}X^{*}$, for any orthonormal matrix $R$, $RZ$ is another solution to the optimization problem 6. Thus, [72] designed an algorithm to find the optimal orthonormal matrix $R$ that converts $X^{*}$ to discrete value in $\{0,1\}^{N\times K}$. From our experiment, [72] is more consistent than K-means and other clustering algorithms, so we stick to using it for our model.
A.3 Data synthesize process
We followed the retina sampling approach described in [14] to achieve foveated imaging. Specifically, each retinal ganglion cell is represented using a Gaussian kernel. The kernel is parameterized by its center, denoted as $\vec{x}_{i}$, and its scalar variance, $\sigma_{i}^{\prime 2}$, i.e. $\mathcal{N}(\vec{x}_{i},\sigma_{i}^{\prime 2}\mathbf{I})$, which is illustrated in Figure 5.A. The response of each cell, denoted as $G[i]$, is computed by the dot product between the pixel value and the corresponding discrete Gaussian kernel. This can be formulated as:
@@ A.3 Data synthesize process
Table 5: Evaluation on foveated CIFAR-10 with varying hyperparameter for density function. For each set of values of $\alpha$ and $\beta$, we perform density-adjusted spectral clustering and run URLOST with the corresponding cluster. The evaluation of each trial is provided in the table.
@@
A.4 Density adjusted spectral clustering on foveated CIFAR10 dataset
We provide further intuition and visualization on why density-adjusted spectral clustering allows the model to learn a better representation of the foveated CIFAR-10 dataset.
As shown in Figure 5, the kernel at the center is much smaller in size than the kernel in the peripheral. This makes the kernel at the center more accurate but smaller, which means it summarizes less information. Spectral clustering with constant density will make each cluster have a similar number of elements in them. Since the kernel in the center is smaller, the cluster in the center will be visually smaller, than the cluster in the peripheral. The effect is shown in Figure 6. Moreover, since we’re upsampling an already low-resolution image (CIFAR-10 image), even though the kernel at the center is more accurate, we’re not getting more information. There, to make sure each cluster has similar information, the clusters in the center need to have more elements than the clusters in the peripheral. In order to make the clusters at the center have more elements, we need to weight the clusters in the center more with the density function. Since the sampling kernels at the center have small eccentricity and are more correlated to their neighbor, increasing $\alpha$ and $\beta$ will make sampling kernels at the center have higher density, which makes the cluster at the center larger. This is why URLOST with density-adjusted spectral clustering performs better than URLOST with constant density spectral clustering, which is shown in Table 5. Meanwhile, setting $\alpha$ and $\beta$ too large will also hurt the model’s performance because it creates clusters that are too unbalanced.
Figure 6: Effect of density adjusted clustering. Eccentricity-based sampling lattice. The center of the sampling lattice has more pixels which means higher resolution compared to the peripheral. (A) Result of density-adjusted spectral clustering ($\alpha=0.5,\beta=2$). Clusters in the center have more elements than clusters in the peripheral. But clusters look more visually similar in size than B. (B) Result of uniform density spectral clustering ($\alpha=0,\beta=0$). Each cluster has a similar number of elements in them but the clusters in the center are much smaller than the clusters in the periphery.
A.5 self-organizing layer learns inverse permutation
@@ A.4 Density adjusted spectral clustering on foveated CIFAR10 dataset
A.6 Training and evaluation details
$\beta$-VAE. $\beta$-VAE was trained for 1000 epochs and 300 epochs on the V1 neura l and TCGA gene expression respectively. We use the Adam optimizer with a learning rate of $0.001$ and a cosine annealing learning rate scheduler. The encoder is composed of a 2-layer MLP with batch normalization and LeakyReLU activation. We use hidden dimensions 2048 and 1024 for V1 and Gene datasets respectively. Then, two linear layers are applied to get the mean and standard for reparameterization. The decoder also has a 2-layer MLP, symmetric to the encoder but using standard ReLU activation and no batch normalization. We tried out different hyperparameters and empirically found that this setting gave the best performance.
MAE. MAE follows the official implementation from the original paper. For CIFAR10, we ran our model for 10,000 epochs. We use Adam optimizer with a learning rate of 0.00015 and a cosine annealing. To fit in our tasks, we use 8 8-layer encoders and 4-layer decoders with hidden dimension 192. The ViT backbone can take different patch sizes and we indicated them accordingly in Table 1. ViT(Pixel) means treating each pixel as a patch, so essentially the patch size is 1. This is also used for the real-world high-dimensional dataset since no concept of patch is defined in the signal space. For V1 neural recording and TCGA gene expression task, we use 4 layers encoder and 2 layers. We use hidden dimension 1380 for 1000 epochs and hidden dimension 384 with 3000 epochs for V1 neural recording and TCGA dataset task. The hidden dimension and the number of epochs we used for MAE are greater than $\beta$-VAE. However, when we use the same parameters on $\beta$-VAE, we did not seem to find a performance gain. Training Transformers usually requires a large number of data. For example, the original transformer on vision is pretrained over 14M images.
URLOST MAE Many hyperparameters are associated with the spectral clustering algorithms. The performance of the model depends on the cluster, which thus depends on these hyperparameters. In general, URLOST is not very sensitive to these hyperparameters. Density parameters (how nodes are normalized) and the number of clusters affects the shape of the clusters, thus, affects the performance of the model to some degree.
Intuition on hyperparameter selection: Both these hyperparameters are related to the size of each cluster, which relates to how difficult the task is. If a cluster is too big, then we need to predict a big missing area on the graph, making the unsupervised learning task too difficult. On the other hand, if the cluster is too small, for example, if the cluster shrinks down to one pixel for an image, then the prediction tasks become too easy. The solution is simply doing a low-pass filter on the graph to fill in the missing pixel. So the model will not learn high-level or semantic representation. Both the number of clusters and density factors are related to the cluster size. Increasing the number of clusters could make each cluster smaller. The density factor essentially defines how much we want to normalize each node on the graph. For an image dataset, clusters should contain semantic sub-structures of the image, like body parts or parts of objects. We use 64 clusters, which results in clusters of size roughly 4x4. In other words, the ratio between the size of clusters (part) and the number of dimensions (whole) is roughly 1:64. For other datasets (gene and V1), we roughly keep this ratio and perform a grid search over these hyperparameters. Similarly, for the density factors, we also perform a grid search centered at the density factor equal to 1. Empirically, URLOST is not very sensitive too these hyperparameters as shown in Figure 7.
We also provide the exact hyperparameters we used for the experiments in this paper: the parameter of URLOST MAE is the same as MAE except for the specific hyper-parameter in the method section. For CIFAR10, we use $K=20$, $\alpha=0.5$ and $\beta=2$. We set the number of clusters to be 64. For V1 neural recording, we use $K=15$, $\alpha=0$ and $\beta=1$. We set the number of clusters to 200. For TCGA dataset, we use $K=10$, $\alpha=0$ and $\beta=1$. We set the number of clusters to 32.
A.7 Effect of spectral clustering hyperparameters on performance
We perform an experiment doing a grid search over each hyperparameter in spectral clustering that affects the performance of URLOST. The grid search experiment is performed over the TCGA gene classification dataset. As shown in Figure 7, the model is not sensitive for most of the hyper-parameters. For a large range of hyperparameters, the model performance stays the same.
Figure 7: Effect of spectral clustering hyperparameters on performance (A) Effect of number of cluster on performance. (B) Effect of normalization parameter for spectral clustering ($\beta$ in Eq.  4) on performance. (C) Effect of numbers of neighbors from $K$ nearest neighbors on performance.
A.8 Topology and Stationarity
We provide a formal definition of topology and stationarity used in the context of our work here:
On topology of the domain of signals. A signal is a function $X(s)$ defined on some domain $S$, where $s\in S$. Let’s focus on $S$. While the definition of topology is a little abstract and relies on open sets, the intuition is that it defines a generalized sense of nearness, or whether two points are close to each other or far apart. A common way to give such a space the property of being near or far away is with an explicit distance function, or a metric. The distance function generates a class of objects—open sets. These spaces (with an associated metric) are called metric spaces. For example, the topology of image space is induced by the L2 norm as the metric function. However, not every space will have a natural metric on it. This is why we need to use the term “topology”.
Topological spaces are generalizations of metric spaces. The definition of topology is any collection of subsets of $S$ that satisfies certain axioms: close under arbitrary union and finite intersection. In the practical setting, the domain $S$ is always discretized as a set of indices, and we can assume that we always deal with a signal defined on a graph [70]. Graphs can be embedded in $\mathbb{R}^{3}$. For a finite graph, the graph topology is just the subspace topology inherited from the usual topology in $\mathbb{R}^{3}$. Under this specific definition, the word “graph,” “adjacency matrix,” and “topology” can be used interchangeably. In this work, we assume we are not provided the topology. Thus, using pairwise mutual information as the adjacency matrix is a way to model the topology. Correlation is another way to model the topology, where we use only second-order pairwise statistics. This is slightly different from the reviewer’s understanding “Topology appears to correspond to ”correlation” between dimensions.” More generally, graphs are topological spaces that are the simplicial 1-complexes or the 1-dimensional CW complexes.
Example with stationarity but no topology: Permuted Cifar10 is an example we had in the manuscript for this criterion (no topology but has stationarity). No topology refers to the fact that $n(i)$ is not defined after random permutation. “Has stationarity” means that the signal is stationary when the topology is recovered. Specifically, we use the mutual information graph to define neighbors $n(i)$. Under this particular definition of $n(i)$, there exists a meaningful invariant statistic $S_{f}(i)$ for the signal. Thus, the signal is stationary. This is because the topology recovered is very similar to the original 2d topology of images before permutation, and the sign before permutation is stationary. However, under an incorrect definition of $n(i)$, the signal is no longer stationary.
Example with no topology but no stationarity: One example would be foveated images simulated using convolutional kernels. One could make a different region of an image have a different resolution while preserving the 2d grid structure of the image by using a Laplace pyramid. As a result, the image is still represented as a 2d matrix, and the definition of $n(i)$ is the same as regular images. However, many invariant statistics before the foveation process are destroyed due to the imbalance resolution at different regions of the image.
On stationarity. In probability theory, a random process is a family of random variables. While the index was originally and frequently interpreted as time, it was generalized to random fields, where the index can be an arbitrary domain. That is, by modern definitions, a random field is a generalization of a stochastic process where the underlying parameter need no longer be real or integer valued ”time” but can instead take values that are multidimensional vectors in $\mathbb{R}^{n}$, points on some manifolds and Lie Groups, or graphs [81; 50; 77; 42; 65; 5]. Let’s consider a real-valued random field $X(s)$ defined on a homogeneous space $S=$$s$ of points $s$ equipped with a transitive transformation group $G=$$g$ of mappings of $S$ into itself, and having the property that the values of the statistical characteristics of this field do not change when elements of $G$ are applied to their arguments. In convention, we call a random field $X(s)$ a strict-sense stationary (or strict-sense homogeneous) random field if for all $n=1,2,\cdots$ and $g\in G$, the finite-dimensional probability distribution of its values at any $n$ points $s_{1},\cdots,s_{n}$ coincides with that of its values at $gs_{1},\cdots,gs_{n}$. We call $X(s)$ field wide-sense stationary (or wide-sense homogeneous) random field if $E|X(s)|^{2}<\inf$ and $EX(s)=EX(gs)$, $EX(s)X(s_{1})=EX(gs)EX(gs_{1})$ for all $s$, $s_{1}$ $\in S$ and $g\in G$. Or wide-sense stationarity means the first and second-order statistics do not change w.r.t. the group action. Further, we can intuitively generalize the wide-sense stationarity to higher-order statistics if we need to define any nth-order stationarity between strict-sense and wide-sense.
In intuitively, stationarity just means the statistics does not change w.r.t. transformations. Let’s take the image as an example, it is a signal $X(s)$ defined on $S=\mathbb{Z}^{2}$. And let’s define the group action as translation. If we use a convolution neural network and learn a set of $L\times L$ filters or use a VIT to learn a set of shared features on $L\times L$ patches. A hidden assumption is that the statistics of our image does not change for $L\times L$ region, or in order words the joint probability $P(X(1,1),\cdots,X(L,L))$ = $P(X(1+a,1+b),\cdots,X(L+a),\cdots,X(L+b))$, where $(a,b)\in$ translation group $T$. While we have initially cited [8], it might be a good idea to provide a more thorough explanation in the appendix, hence, we have included another five references to provide a more comprehensive point of view.
A.9 Computational cost
Computation cost. The computational cost and scalability of URLOST are nearly the same as MAE. The computational cost of MAE is analyzed in the original MAE paper. Its efficient design of MAE speeds up the training and evaluation time around 4x compared to a standard Transformer-based auto-encoder. The additional elements in URLOST are simply clustering and self-organizing layers. This effect is shown in figure 7, despite the encoder used in the experiment being 3 times as the decoder, the runtime of the encoder is still faster than the decoder. Moreover, the operation of clustering indexing and self organizing layer is very small relative to the overall runtime of the model. The mutual information and the clustering results are precomputed based on the dataset and thus do not introduce more cost for the model training. During training and inference, clustering is just a one-step indexing process that does not create significant overhead in additional to MAE. For all experiments performed in this paper, since clustering and mutual information are computed and saved for all datasets at once, their computational cost is negligible with respect to the training time of MAE. However, The runtime of these two procedure could be a problem for extremely high dimensional data.
With the current implementation, we perform a scalability benchmark as suggested by the reviewer. We resize an image dataset to different resolutions. At each resolution, we compute mutual information and clustering for a different number of pixels (dimensions), and we provide a plot of running time with respect to the number of pixels in Figure 8. Both computational cost of computing pairwise mutual information and clustering both scale quadratically with the number of dimensions of the signal. It takes aroudn 130 minutes to calculate mutual information for data with 10,000 dimensions. Clustering takes much shorter time (less than a minute).
Figure 8: Runtime for clustering and computation mutual information. A. Benchmark of runtime for clustering for signal with different dimensions. B. Benchmark of runtime for computing pairwise mutual information with different dimensions. Both run time scale quadratically as the dimension of the signal.
Amount of parameters. The self-organizing layer introduces additional parameters compared to the original MAE model. We believe additional parameters are a reasonable trade-off for not knowing topology and stationarity, which are important inductive biases in the data. However, it doesn’t increase the parameters drastically, we provide the ratio of the additional parameters in table 6.
Table 6: Ratio between Parameters of Self-Organizing Layers and Parameters of the Entire Model. We report the model used for each dataset.
Dataset Permuted Cifar10 TCGA V1
Ratio 6.7% 15.6% 5.9%
Table 7: Runtime of each module during the forward operation. We report the runtime of each module during inference for a single example. We use the model trained on permuted CIFAR10 dataset. The experiment is performed using a single RTX 2080TI and is averaged over 500 trials.
Module Cluster Indexing Self Organizing Layer Encoder Decoder
Runtime 0.27 ms 1.11 ms 17.5 ms 20.4 ms
A.10 Scalability
Scalability is not the primary focus of this work. In general, unsupervised representation learning benefits from scaling to a larger amount of data, and the MAE model, which our method is built upon, demonstrates a level of scaling [34]. We believe URLOST inherits this potential for the fact that it shows stronger advantages on more challenging tasks, as seen with CIFAR-10 compared to simpler benchmarks. This suggests that scaling to larger and more complex datasets is a promising direction. Our study explores the novel settings of unsupervised representation learning without knowing the data topology and stationarity, demonstrating the effectiveness of our method across diverse data modalities with significant differences. This generalization is our key and positive contribution, highlighting the potential of our approach to inspire future work in this area.
A.11 Noise robustness of mutual information recovered topology
We tested the robustness of the topology recovery algorithm used in URLOST. Specifically, we add gaussian noise to CIFAR10 to generate data with a low signal-to-noise ratio.
We found out that while decreasing signal-to-noise ratio will lower the overall mutual information between pairs of pixels, the relationship between mutual information between pixels and image topology (distance between two pixels) does not change. We perform the following experiments: we add iid gaussian noise to each pixel with a fixed noise level, then we measure the mutual information between pairs of pixels as $MI((i,j),(k,l))$. For visualization, we aggregate all pairs by the distance between them and average their mutual information, i.e. let $A(a)={(i,j),(k,l)s.t.\sqrt{(i-k)^{2}+(j-l)^{2}}}$. We have $MI(a)=\frac{1}{|A(a)|}\sum_{i,j,k,l\in A(a)}MI((i,j),(k,l))$. $MI(a)$ denotes the mutual information between pixels of different distances.
As shown in Figure 9 below, although the overall mutual information decreases a lot as the noise level increases, the order of MI(a) is still the same. In other words, the mutual information is still the biggest when the two pixels are nearby and decays when the mutual information decreases. This implies the topology recovered by pairwise mutual information is robust even when the signal has a lower signal-to-noise ratio. It only breaks when you add too much noise ($\sigma=0.9$). For this amount of noise, the image is not distinguishable.
Figure 9: The effect of additive noise on pairwise mutual information. We plot averaged mutual information between pairs of pixels. We aggregate all pairs with the same distance. Each curve represents a specific noise level. Although the curve becomes lower as noise increases, the decaying features of the curve remain the same.
A.12 Self-Organizing Map (SOM)
Kohonen in 1982 presents a fascinating thought experiment [43]: “Suppose you woke up one day to find someone rewired your optic nerve (or you have been implanted with a prosthetic retina). The signals from the retina to the brain are intact, but the wires are all mixed up, projecting to the wrong places. Can the brain learn to “descramble” the image?” Kohonen proposes the “self-organizing map” (SOM) algorithm to address this problem. SOM is an unsupervised machine learning technique used to produce a low-dimensional (typically two-dimensional) representation of a higher-dimensional dataset while preserving the topological structure of the data. In the task of “descrambling” pixels, SOM maps each pixel into a 2D index set by leveraging the mutual information between pixels at different locations. The parameters of SOM include a set of neurons, where each neuron has a location index. Let $\mathbf{W}_{v}$ denote the weight vector of the $v$th neuron and $\mathbf{D}$ the input vector. $\theta(u,v,s)$ is the interaction function that determines the influence between the $u$th and $v$th neurons based on their distance. The update equation for SOM is given by:
$\mathbf{W}_{v}(s+1)=\mathbf{W}_{v}(s)+\theta(u,v,s)\cdot\alpha(s)\cdot\left(\mathbf{D}(t)-\mathbf{W}_{v}(s)\right)$
In the task of “descrambling” pixels, the input consists of pixel values. The indices will be two-dimensional, such as $s=(i,j)$, and $\mathbf{W}_{v}(s)$ will learn to represent a set of “descrambled” pixels, where $s$ represents the correct indices of the pixels. In other words, the index set defines the correct topology of the image as a 2D grid, and SOM maps the pixels onto this 2D topology by leveraging the mutual information between pixels at different locations. Other methods such as [67] use manifold learning to uncovered topology 2-d topology from natural images. However, when the intrinsic dimension of the manifold is larger than 2d, it is difficult to integrate the ”uncovered” topology with state-of-the-art self-supervised learning algorithms.
A.13 Visualizing the weight of self-organizing